20% of items produced from a factory are defective. What is the probability that in a sample of 5 chosen at random, none is defective, one is defective and p (1<x<4)?


1
Expert's answer
2022-02-14T16:27:19-0500

Let "X=" the number of defective items: "X\sim Bin(n, p)."

Given "n=5, p=0.2"

i)


"P(X=0)=\dbinom{5}{0}(0.2)^0(1-0.2)^{5-0}=0.32768"

ii)


"P(X=1)=\dbinom{5}{1}(0.2)^1(1-0.2)^{5-1}=0.4096"

iii)


"P(1<X<4)=P(X=2)+(X=3)"

"=\dbinom{5}{2}(0.2)^2(1-0.2)^{5-2}+\dbinom{5}{3}(0.2)^3(1-0.2)^{5-3}"

"=0.2048+0.0512=0.256"